A company manufactures and sells bar stools. Manufacturing costs consist of a fixed cost of R 25000 per month and a variable cost of R350.00 per unit. Each unit is sold at R600.00. determine the linear profit equation and what is the break even level. assignment experts
1
Expert's answer
2018-05-13T07:24:08-0400
Linear profit equation is
π=600q-(25000+350q)=250q-25000
Break even level is
q=25000/250=100